# Non-Gaussian state preparation and enhancement using weak-value amplification

### v1

@@ III Non-Gaussian sates generation
Figure 2: Fidelity functions between the signal output state $|\Phi^{\prime}\rangle$ and the ideal SPAC state (solid curve), as well as the coherent state $|\phi\rangle$ (dashed curve). Here, $\epsilon=0.001$ and $g=0.01$.
We observe that the state $|\Phi^{\prime}\rangle$ is a superposition of a coherent state $|\beta\rangle$ and a SPAC state $|\phi_{1}\rangle=b^{\dagger}|\beta\rangle$. When the weak value $\langle n_{a}\rangle_{w}$ is large, the contribution of the SPAC component $|\phi_{1}\rangle$ dominates over the coherent component. In general, the generated state can be regarded as a typical example of a hybrid coherent state of the form $\sqrt{\varepsilon}|\beta\rangle+\sqrt{1-\varepsilon}|\phi_{1}\rangle$, with $0\leq\varepsilon\leq 1$ (Turek et al., 2023). Such states exhibit enhanced nonclassical features compared to the SPAC state alone because of their coherent-state admixture in specific parameter regimes.
@@
To further characterize the output state, we examine its Wigner function, which provides a complete description of the phase-space distribution of the state. For a quantum state with density operator $\rho$, the Wigner function is obtained via the Fourier transform of the characteristic function $C_{w}(\lambda)=Tr\left[\rho D(\lambda)\right]$. In this study, we used the Python package QuTiP (Johansson et al., 2012, 2013) to numerically compute the Wigner functions of the relevant quantum states.
Figure 3: Wigner function of the output state of the generated signal $|\Phi^{\prime}\rangle$. (a) $\beta=0$; (b) $\beta=0.5$; (c) $\beta=1$. Other parameters are the same as those used in Fig. 2.
To visualize the evolution of the phase-space distribution, Fig.3 shows the Wigner functions of the output state $|\Phi^{\prime}\rangle$ for different values of $\beta$. We fixed the parameters as $\epsilon=0.001$ and $g=0.01$, and considered $\beta=0$, $0.5$, and $1$. Figure3(a) corresponds to the input coherent state, exhibiting the expected Gaussian distribution displaced in the phase space. In Fig.3(b), for $\beta=0.5$, the Wigner function of the output state shows significant negativity near the origin, along with squeezing along the $x$-quadrature. This indicates strong nonclassicality and a marked deviation from the initial Gaussian character. In Fig.3(c), as $\beta$ increases further, the negativity gradually diminishes, and the Wigner function approaches a Gaussian profile again (not shown). This trend is consistent with the behavior of the fidelity functions in the large-$\beta$ regime [see Fig. 2]. These numerical results demonstrate that our protocol effectively transforms the initially Gaussian coherent state into a nG state with tunable nonclassicality, depending on the input amplitude $\beta$.
@@
Because the SPAC state lacks a vacuum component and exhibits a high probability of single-photon detection, it has found various applications in quantum information processing, including quantum key distribution (QKD) (Wang et al., 2014). In practical QKD implementations, sources with sub-Poissonian photon-number statistics and a dominant single-photon component are highly desirable. Based on the above analysis, our generated state $|\Phi^{\prime}\rangle$ satisfies both of these criteria, making it a promising candidate for QKD applications. Therefore, $|\Phi^{\prime}\rangle$ could offer improved performance compared to other commonly used sources, such as weak coherent states (Zhao et al., 2006; Rosenberg et al., 2007) and heralded single-photon sources (Wang et al., 2008, 2007).
Figure 4: Second-order correlation function $g^{(2)}(0)$. The blue solid curve corresponds to the generated signal output state $|\Phi^{\prime}\rangle$, the red and green dashed curves represent the coherent state and ideal SPAC state cases, respectively. Other parameters are the same as those used in Fig. 2.
III.2 Squeezed photon number state
@@
In the previous subsection, we considered a coherent state as the signal input and demonstrated that the resulting output states exhibited nG characteristics. In this subsection, we consider another typical example, a single-mode SV state. The SV state is defined as $|r\rangle=S(r)|0\rangle$, where the squeezing operator is given by $S(r)=\exp\left[\frac{1}{2}(r^{\ast}b^{2}-rb^{\dagger 2})\right]$, and $r$ is the squeezing parameter. For $r>0$, the SV state is squeezed in the position quadrature $x=\frac{1}{\sqrt{2}}(b+b^{\dagger})$ and anti-squeezed in the momentum quadrature $p=\frac{-i}{\sqrt{2}}(b-b^{\dagger})$; the reverse holds for $r<0$. The squeezed vacuum state is a fundamental resource in quantum optics (Braunstein, 2005), with substantial interest owing to its distinctive properties and wide-ranging applications in quantum sensing (Polzik et al., 1992; et al., 2011), quantum communication (Braunstein and van Loock, 2005), and others (see, e.g., Refs. (Vahlbruch et al., 2016; Andersen et al., 2016; Schnabel, 2017)). In addition, the photon-added and photon-subtracted SV states have been studied extensively (Parigi et al., 2007). In general, generating multiphoton-added or multiphoton-subtracted SV states suffers from very low success probabilities due to measurement imperfections and operational inefficiencies. From the implementation perspective, photon subtraction is typically easier than photon addition. Therefore, the preparation of multiphoton-added SV states requires advanced techniques and higher efficiency. Herein, we propose a simple method to generate two-photon-added squeezed vacuum (TPASV) states without requiring actual photon-addition operations. In the schematic of our setup, if the initial signal (pointer) state is prepared as an SV state, the final signal output state after the measurement process can be expressed as
@@
$\mu=\frac{i\frac{g}{\sqrt{2}}\left(1+\frac{1}{\epsilon}\right)sinh\left(r\right)cosh\left(r\right)}{1-i\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)sinh^{2}\left(r\right)}.$ (16)
@@
$|\Psi\rangle=\eta\left[|r\rangle-ig\langle n_{a}\rangle_{w}|\psi_{b^{\dagger 2}}\rangle\right],$ (17)
where
$|\psi_{b^{\dagger 2}}\rangle=\mathcal{N}_{a^{\dagger 2}}S\left(r\right)\left[|0\rangle-\sqrt{2}\left(\tanh r\right)^{-1}|2\rangle\right]$ (18)
is the TPASV state with normalization constant $\mathcal{N}_{a^{\dagger 2}}=[1+2(\tanh r)^{-2}]^{-1/2}$. The normalization constant $\eta$ for the state $|\Psi\rangle$ is given by
@@
Figure 5: Wigner functions (contour plots) and quadrature distributions for the generated state $|\Psi\rangle$ with different squeezing parameters $r$. (a) and (c) show the Wigner functions of the signal input SV states with squeezing parameters $r=-0.7$ and $r=0.2$, respectively. The corresponding Wigner functions of the output signal states $|\Psi\rangle$ are shown in (b) and (d). Other parameters are the same as those used in Fig. 2.
@@
$b^{\dagger 2}S\left(r\right)|0\rangle=b^{\dagger}bS\left(r\right)|0\rangle.$ (20)
This relation can be verified using the following operator identities:
@@
From Eqs. (21a) and (21b), we see that the squeezed single-photon state $S(r)|1\rangle$ can be obtained by adding or subtracting a single photon to or from an SV state. As confirmed in previous studies (Lund et al., 2004; Ourjoumtsev et al., 2006), the SV state and squeezed single-photon state are good approximations of small-amplitude even and odd SC states, respectively. The SC states are defined as
@@
As noted above, the generated state $|\Psi\rangle$ takes the form $c_{0}|0\rangle+c_{2}|2\rangle$, with $c_{0},c_{2}\in\mathbb{C}$. That is, it is a squeezed superposition of the vacuum and two-photon Fock states. Such a state provides an excellent approximation of a squeezed even SC state with large amplitude $\alpha$ (Marek et al., 2008), which is defined as
@@
where the normalization constant is $\mathcal{N}_{sscs}=\left(2+2e^{-2\alpha^{2}}\right)^{-1/2}$. This approximation is valid because the squeezing operator $S(r^{\prime})$, with an appropriately chosen squeezing parameter $r^{\prime}$, effectively pulls apart the two peaks of the cat state. To explain this effect more clearly, in Fig. 5, we plotted the Winger functions of our input SV and the corresponding signal output state $|\Psi\rangle$ for different squeezing parameter $r$. Using our protocol, the input Gaussian-type SV state transformed to the nG state characterized by Winger negativity, and squeezing occurred in both $x$ and $p$ quadratures [ see Figs. 5 (b) and (d) ]. The interference fringes and quadrature distribution features of the state $|\Psi\rangle$ showed in Fig. 5 is very similar to the squeezed even SC state which is the basis of GKP states (Gottesman et al., 2001; Hastrup and Andersen, 2022). It also indicates that the two peaks of the generated state are separated along $x$ or $p$ directions depending on the value of the squeezing parameter $r$. In the following, we further analyze the similarity between our generated state $|\Psi\rangle$ and the ideal large-amplitude squeezed even SC state using fidelity and phase-space distributions. In the following, we assume that both squeezing parameters $r$ and $r^{\prime}$ are real.
@@
In Fig. 6, we present the optimal fidelity between $|\Psi\rangle$ and $|\Psi_{sscs}\rangle$ as a function of the coherent amplitude $\alpha$ for different squeezing parameter $r$. In these plots, for each value of $\alpha$, we optimize the squeezing parameter $r^{\prime}$ of the ideal squeezed even SC state to maximize its overlap with the generated state $|\Psi\rangle$. The cases $r=-0.7$ and $r=-2$ are shown in panels (a) and (b), respectively. The corresponding optimized values of $r^{\prime}$ as a function of $\alpha$ are plotted in panels (c) and (d), respectively.
In Figs.6(a) and 6(b), the blue solid curves represent the optimal fidelity between $|\Psi\rangle$ and $|\Psi_{sscs}\rangle$, while the green dashed curves show the fidelity between $|\Psi\rangle$ and the ideal even SC state (without squeezing). As shown in Fig.6(a) for $r=-0.7$, the optimal fidelity reaches as high as $F_{2}=0.97$ at $\alpha=1.91$ and $r^{\prime}=-0.14$, indicating a high degree of similarity between the generated state and the ideal squeezed even SC state. As shown in Fig. 6(b), when $r=-2$, the output signal state $|\Psi\rangle$ provides an excellent approximation to the squeezed even SC state, achieving a very high fidelity of $F_{2}=0.994$ at $\alpha=2.07$ and $r^{\prime}=-0.192$. This result indicates that, in this case, the output signal state $|\Psi\rangle$ is extremely well approximated by the ideal squeezed even SC state.
Figure 6: (a) and (b) show the optimal fidelities between the output signal state $|\Psi\rangle$ and the target squeezed even SC state (solid curves), as well as the ideal even SC state (dashed curves), for different system parameters. (c) and (d) show the corresponding squeezing parameter $r^{\prime}$ of the target squeezed even SC state for which the fidelities in (a) and (b) are optimized. The squeezing parameters of the initial SV states are $r=-0.7$ and $r=-2$ for (a) and (b), respectively. The other parameters are the same as those used in Fig. 2.
@@
Figure 7(a) shows the Wigner function of $|\Psi\rangle$ for $r=-0.7$, while Fig.7(b) displays the Wigner function of the squeezed even SC state with $r^{\prime}=-0.14$ and $\alpha=2$. For these parameters, the fidelity between the two states is approximately $F_{2}\approx 0.97$. It also further confirmed the high similarity of our generated output signal state $|\Psi\rangle$ to the $|\Psi_{sscs}\rangle$ with large amplitude $\alpha$. Fig.7 also illustrates the effects of our protocol on the initially prepared Gaussian SV state, which transforms it into a nG state.
Figure 7: (a) Wigner function of the generated state $|\Psi\rangle$ with $r=-0.7$. (b) Wigner function of the ideal squeezed even SC state with $r^{\prime}=-0.14$ and $\alpha=2$. Other parameters are the same as those used in Fig. 2. For these parameters, the fidelity calculated using the Wigner functions of the generated and ideal squeezed even SC states is $F_{2}\approx 0.97$.
As mentioned in the Introduction, the generation of SC states with high fidelity ($F\geq 0.99$) while maintaining large coherent amplitudes ($\alpha\geq 2$) is a crucial task. Although our scheme does not directly generate large-amplitude SC states, the output signal state achieves exceptionally high fidelity with a squeezed even SC state possessing a large amplitude $\alpha$. Therefore, our approach could be applied to problems where large-amplitude SC states are required, offering potentially better performance.
IV Enlargement of Schrödinger cat states
@@
where $\kappa^{\prime}=g\langle n\rangle_{w}$ and $\mathcal{N}_{\pm}^{\prime}$ is the normalization constant, given by
@@
Figure 8: (a) and (c) show the optimal fidelities between the output signal states $|\Psi_{\pm}\rangle$ and the ideal SC states $|SCS_{\pm}(\alpha^{\prime})\rangle$, plotted as a function of the coherent state amplitude $\alpha$. (b) and (d) show the corresponding enlargement coefficients $\eta_{-}=\alpha^{\prime}/\alpha$ and $\eta_{+}=\alpha^{\prime}/\alpha$, for the odd and even SC states, respectively, for which the fidelities in (a) and (c) are optimized. The other parameters are the same as those used in Fig. 2.
In Fig. 8, we show the optimal fidelity functions $F_{+}$ and $F_{-}$ between our generated output states $|\Psi_{+}\rangle$ and $|\Psi_{-}\rangle$ and the ideal even and odd SC states, respectively, as functions of coherent state amplitude $\alpha$. The corresponding optimized amplitudes $\alpha^{\prime}$ of the ideal SC states are used for the maximal overlap. As shown in Fig. 8(a), the generated state $|\Psi_{-}\rangle$ provides an excellent approximation to the ideal odd SC state, with fidelity $F_{-}\geq 0.965$ across a wide range of amplitudes, $\alpha$. Notably, in the regions $0.80<\alpha<$ 0.8 and $\alpha\geq 2.45$, the fidelity exceeds 0.99. For the even SC state input case [Fig. 8(c)], the generated state $|\Psi_{+}\rangle$ initially shows noticeable deviation from the ideal even SC state in the small-amplitude regime ($0.290<\alpha<$ 0.29). However, as $\alpha$ increases beyond 0.29, the fidelity $F_{+}$ improves significantly, surpassing $0.9$ for all $\alpha>1.24$ regions. Thus, for $\alpha>1.24$, $|\Psi_{+}\rangle$ also can be considered an excellent approximation to the ideal even SC state. Moreover, in high-fidelity regions, we observed that the optimized coherent state amplitudes $\alpha^{\prime}$ of the ideal SC states are consistently larger than the corresponding values of $\alpha$ for our generated states $|\Psi_{\pm}\rangle$ [see Figs. 8(b) and 8(d)]. This indicates that our protocol effectively enlarges the SC states. This enlargement effect can be further confirmed by comparing the Wigner functions of the input ideal SC states and the generated output states $|\Psi_{\pm}\rangle$.
@@
In the small-amplitude regime ($\alpha\leq 1.2$), the odd SC state is well approximated by the squeezed single-photon state $S(r)|1\rangle$, with a fidelity exceeding $0.99$ (Zavatta et al., 2004; Ourjoumtsev et al., 2006). The corresponding optimized squeezing parameter that maximizes the fidelity for a given $\alpha$ is given by $r_{opt}(\alpha)=\ln\left(\sqrt{\frac{2\alpha^{2}}{3}+\frac{1}{3}\sqrt{9+4\alpha^{2}}}\right)$. However, for larger amplitudes ($\alpha\geq 2$), the fidelity rapidly decreased below $87.8\%$. In contrast, we find that our protocol can generate odd SC states with high fidelity, even for large amplitudes $\alpha\geq 2$. In Fig. 9, we plot the Wigner functions of the odd SC input states for different values of $\alpha$, along with the corresponding output states $|\Psi_{-}\rangle$, using the optimized amplitude $\alpha^{\prime}$ which maximizes the fidelity between ||||$\Psi_{-}\rangle$ and the ideal odd SC target state $|SCS_{-}(\alpha)\rangle$. As shown, for each input odd SC state, our protocol yields an excellent approximation to a larger-amplitude odd SC state encoded in the output state $|\Psi_{-}\rangle$. For example, when the input amplitudes are $\alpha=0.5$, $1.52$, and $2.0$, the corresponding enlarged SC states, characterized by the output amplitudes $\alpha^{\prime}=0.85$, $2.00$, and $2.41$, are obtained with high fidelities: $F_{-}=0.999$, $0.968$, and $0.982$, respectively. The amplitude enlargement ratio $\eta_{-}=\alpha^{\prime}/\alpha$ is plotted in Fig. 8(b). The curve shows that $\eta_{-}>1$ across a wide range of $\alpha$, indicating that our protocol effectively amplifies the amplitude of a given odd SC state while maintaining high fidelity.

### v2

@@ A Versatile Protocol via Weak Measurement
Figure 2: Fidelity between the generated output state and reference states. Fidelity functions between the signal output state $|\Phi^{\prime}\rangle$ and the ideal SPAC state (solid curve), as well as the coherent state $|\phi\rangle$ (dashed curve). Here, $\epsilon=0.001$ and $g=0.01$.
We observe that the state $|\Phi^{\prime}\rangle$ is a superposition of a coherent state $|\beta\rangle$ and a SPAC state $|\phi_{1}\rangle=b^{\dagger}|\beta\rangle$. When the weak value $\langle n_{a}\rangle_{w}$ is large, the contribution of the SPAC component $|\phi_{1}\rangle$ dominates over the coherent component. In general, the generated state can be regarded as a typical example of a hybrid coherent states of the form $\sqrt{\varepsilon}|\beta\rangle+\sqrt{1-\varepsilon}|\phi_{1}\rangle$, with $0\leq\varepsilon\leq 1$ (Turek_2023). Such states exhibit enhanced nonclassical features compared to the SPAC state alone due to the admixture of a coherent-state in specific parameter regimes.
The conventional method for generating SPAC states, as demonstrated by Zavatta et al. (2004), employs conditional parametric down-conversion in a $\chi^{(2)}$ nonlinear crystal, where the detection of an idler photon heralds the addition of a photon to a coherent state in the signal mode. While effective, this approach is inherently specialized for producing SPAC states and is constrained by weak second-order nonlinearity and finite detection efficiency, resulting in a low success probability. In contrast, our WVA based protocol, utilizing a weak $\chi^{(3)}$ cross-Kerr interaction, offers a versatile and tunable alternative. Although our scheme also features a low probability due to the required postselection, it is not limited to a single state type. By adjusting system parameters ($\beta$, $\epsilon$, g), our single experimental setup can be tuned to prepare not only SPAC states but also hybrid coherent states (Turek_2023) and other nG states, depending on the initial signal input.
@@
To further characterize the output state, we examine its Wigner function, which provides a complete description of the phase-space distribution of the state. For a quantum state with density operator $\rho$, the Wigner function is obtained via the Fourier transform of the characteristic function $C_{w}(\lambda)=Tr\left[\rho D(\lambda)\right]$. In this study, we used the Python package QuTiP (JOHANSSON20121760; JOHANSSON20131234) to numerically compute the Wigner functions of the relevant quantum states.
Figure 3: Wigner function of the output state of the generated signal $|\Phi^{\prime}\rangle$. a $\beta=0$; b $\beta=0.5$; c $\beta=1$. Other parameters are the same as those used in Fig. 2.
To visualize the evolution of the phase-space distribution, Fig. 3 shows the Wigner functions of the output state $|\Phi^{\prime}\rangle$ for different values of $\beta$. We fixed the parameters as $\epsilon=0.001$ and $g=0.01$, and considered $\beta=0$, $0.5$, and $1$. Figure 3a corresponds to the input coherent state, exhibiting the expected Gaussian distribution displaced in the phase space. In Fig. 3b, for $\beta=0.5$, the Wigner function of the output state shows significant negativity near the origin, along with squeezing in the $x$-quadrature. This indicates strong nonclassicality and a marked deviation from the initial Gaussian character. In Fig. 3c, as $\beta$ increases further, the negativity gradually diminishes, and the Wigner function approaches a Gaussian profile again (not shown). This trend is consistent with the behavior of the fidelity functions in the large-$\beta$ regime [see Fig. 2]. These numerical results demonstrate that our protocol effectively transforms the initially Gaussian coherent state into a nG state with tunable nonclassicality, depending on the input amplitude $\beta$.
Figure 4: Second-order correlation function $g^{(2)}(0)$. The blue solid curve corresponds to the generated signal output state $|\Phi^{\prime}\rangle$, while the red and green dashed curves represent the coherent state and ideal SPAC state, respectively. Other parameters are the same as those used in Fig. 2.
@@
Since the SPAC state lacks a vacuum component and exhibits a high probability of single-photon detection, it has found various applications in quantum information processing, including quantum key distribution (QKD) (PhysRevA.90.062315). In practical QKD implementations, sources with sub-Poissonian photon-number statistics and a dominant single-photon component are highly desirable. Based on the above analysis, our generated state $|\Phi^{\prime}\rangle$ satisfies both of these criteria, making it a promising candidate for QKD applications. Therefore, $|\Phi^{\prime}\rangle$ could offer improved performance compared to other commonly used sources, such as weak coherent states (PhysRevLett.96.070502; PhysRevLett.98.010503) and heralded single-photon sources (PhysRevLett.100.090501; PhysRevA.75.012312).
The success probability of our signal output state $|\Phi^{\prime}\rangle$ is equal to $\epsilon^{2}|\mathcal{N}_{c}|^{-2}$. By numerically analyzing the successful postselection probability of our state, we notice that it is on the order of $10^{-5}$ for single-trial with high fidelity ($F_{1}>0.995$). However, this low yield characteristic is shared by many advanced quantum state generation schemes, including the pioneering work of Zavatta et al. (2004) and Ourjoumtsev et al. (2006). The key is that our protocol is heralded. The successful generation of $|\Phi^{\prime}\rangle$ is signaled by a click at the idler detector ($D_{1}$). This means the experiment can be run continuously and at high repetition rates (e.g., with a pulsed laser). For instance, even with a success probability as low as $10^{-5}$, a 1010MHzMHz to 11GHzGHz repetition rate (Ma_2015; XUE2024107048; CHEN2025111703) yields about $10^{2}\thicksim 10^{4}$ heralded states per second. For many proof-of-principle quantum information protocols including QKD, this rate is entirely practical.
@@
Having demonstrated nG state generation from coherent inputs, we now apply our protocol to a SV input. We will show that this process generates a state equivalent to a squeezed photon number state. In the previous subsection, we considered a coherent state as the signal input and demonstrated that the resulting output states exhibited nG characteristics. In this subsection, we consider another typical example, a single-mode SV state. The SV state is defined as $|r\rangle=S(r)|0\rangle$, where the squeezing operator is given by $S(r)=\exp\left[\frac{r}{2}(b^{2}-b^{\dagger 2})\right]$, and $r$ is the squeezing parameter. For $r>0$, the SV state is squeezed along the position quadrature $x=\frac{1}{\sqrt{2}}(b+b^{\dagger})$ and anti-squeezed in the momentum quadrature $p=\frac{-i}{\sqrt{2}}(b-b^{\dagger})$; the reverse holds for $r<0$. The squeezed vacuum state is a fundamental resource in quantum optics (PhysRevA.71.055801), with substantial interest owing to its distinctive properties and wide-ranging applications in quantum sensing (PhysRevLett.68.3020; 2011), quantum communication (RevModPhys.77.513), and others [see, e.g., Refs. (PhysRevLett.117.110801; Andersen2016; SCHNABEL20171)]. In addition, the photon-added and photon-subtracted SV states have been studied extensively (2007PV). In general, generating multiphoton-added or multiphoton-subtracted SV states suffers from very low success probabilities due to measurement imperfections and operational inefficiencies. From an implementation perspective, photon subtraction is typically easier than photon addition. Therefore, the preparation of multiphoton-added SV states requires advanced techniques and higher efficiency. Herein, we propose a simple method to generate two-photon-added squeezed vacuum (TPASV) states without requiring actual photon-addition operations. In the schematic of our setup, if the initial signal (pointer) state is prepared as an SV state, the final signal output state after the measurement process can be expressed as
@@
$\mu=\frac{i\frac{g}{2\sqrt{2}}\left(1+\frac{1}{\epsilon}\right)sinh\left(2r\right)}{1-i\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)sinh^{2}\left(r\right)}.$ (16)
@@
$|\Psi\rangle$ $=\eta\left[|r\rangle-ig\sinh^{2}r\langle n_{a}\rangle_{w}\mathcal{N}_{b^{\dagger 2}}^{-1}|\psi_{b^{\dagger 2}}\rangle\right],$ (17)
where the normalized expression of the TPASV state $b^{\dagger 2}|r\rangle$ is defined by
$|\psi_{b^{\dagger 2}}\rangle=\mathcal{N}_{b^{\dagger 2}}S\left(r\right)\left[|0\rangle-\sqrt{2}\left(\tanh r\right)^{-1}|2\rangle\right]$ (18)
with normalization constant $\mathcal{N}_{b^{\dagger 2}}=[1+2(\tanh r)^{-2}]^{-1/2}$. The normalization constant $\eta$ for the state $|\Psi\rangle$ is given by
@@
For the parameters considered herein, the overall success probability of generating the state $|\Psi\rangle$ is approximately $10^{-4}$, as determined by the expression $\epsilon^{2}|\mathcal{\eta|}^{-2}$.
Figure 5: Wigner functions (contour plots) and quadrature distributions for the generated state $|\Psi\rangle$. a and c show the Wigner functions of the signal input SV states with squeezing parameters r=0.7r=-0.7 and $r=0.2$, respectively. The corresponding Wigner functions of the output signal states $|\Psi\rangle$ are shown in b and d. Other parameters are the same as those used in Fig. 2.
@@
$b^{\dagger 2}|r\rangle=-\coth r\ b^{\dagger}b|r\rangle.$ (20)
It follows that the normalized TPASV state is identical to the normalized photon-subtracted-and-added squeezed state: $|\psi_{b^{\dagger 2}}\rangle=|\psi_{b^{\dagger}b}\rangle$. This relation can be verified using the following operator identities:
@@
From Eqs. (21a) and (21b), we see that the squeezed single-photon state $S(r)|1\rangle$ can be obtained by adding or subtracting a single photon to or from an SV state. As confirmed in previous studies (PhysRevA.70.020101; 2006), the SV state and squeezed single-photon state are good approximations of small-amplitude even and odd SC states, respectively. The SC states are defined as
@@
As noted above, the generated state $|\Psi\rangle$ takes the form $c_{0}|0\rangle+c_{2}|2\rangle$, with $c_{0},c_{2}\in\mathbb{C}$. That is, it is a squeezed superposition of the vacuum and two-photon Fock states. Such a state provides an excellent approximation of a squeezed even SC state with large amplitude $\alpha$ (PhysRevA.78.063811), which is defined as
@@
where the normalization constant is $\mathcal{N}_{sscs}=\left(2+2e^{-2\alpha^{2}}\right)^{-1/2}$. This approximation is valid because the squeezing operator $S(r^{\prime})$, with an appropriately chosen squeezing parameter $r^{\prime}$, effectively pulls apart the two peaks of the cat state. To explain this effect more clearly, in Fig. 5, we plotted the Wigner functions of our input SV and the corresponding signal output state $|\Psi\rangle$ for different squeezing parameters $r$. Using our protocol, the input Gaussian-type SV state transformed to the nG state characterized by Wigner negativity, and squeezing occurred along both $x$ and $p$ quadratures [ see Figs. 5 b and d ]. The interference fringes and quadrature distribution features of the state $|\Psi\rangle$ showed in Fig. 5 is very similar to the squeezed even SC state which is the basis of GKP states (PhysRevA.64.012310; PhysRevLett.128.170503). It also indicates that the two peaks of the generated state are separated along the $x$ or $p$ direction depending on the value of the squeezing parameter $r$. In the following, we further analyze the similarity between our generated state $|\Psi\rangle$ and the ideal large-amplitude squeezed even SC state using fidelity and phase-space distributions. In the following, we assume that both squeezing parameters $r$ and $r^{\prime}$ are real.
@@
In Fig. 6, we present the optimal fidelity between $|\Psi\rangle$ and $|\Psi_{sscs}\rangle$ as a function of the coherent amplitude $\alpha$ for different squeezing parameter $r$. In these plots, for each value of $\alpha$, we optimize the squeezing parameter $r^{\prime}$ of the ideal squeezed even SC state to maximize its overlap with the generated state $|\Psi\rangle$. The cases r=0.7r=-0.7 and r=2r=-2 are shown in panels a and b, respectively. The corresponding optimized values of $r^{\prime}$ as a function of $\alpha$ are plotted in panels (c) and (d), respectively.
In Figs. 6a and 6b, the blue solid curves represent the optimal fidelity between $|\Psi\rangle$ and $|\Psi_{sscs}\rangle$, while the green dashed curves show the fidelity between $|\Psi\rangle$ and the ideal even SC state (without squeezing). As shown in Fig. 6a for r=0.7r=-0.7, the optimal fidelity reaches as high as $F_{2}=0.97$ at $\alpha=1.91$ and $r^{\prime}=-0.14$, indicating a high degree of similarity between the generated state and the ideal squeezed even SC state. As shown in Fig. 6b, when r=2r=-2, the output signal state $|\Psi\rangle$ provides an excellent approximation to the squeezed even SC state, achieving a very high fidelity of $F_{2}=0.994$ at $\alpha=2.07$ and $r^{\prime}=-0.192$. This result indicates that, in this case, the output signal state $|\Psi\rangle$ is extremely well approximated by the ideal squeezed even SC state.
Figure 6: Optimal fidelities and squeezing parameters. a and b show the optimal fidelities between the output signal state $|\Psi\rangle$ and the target squeezed even SC state (solid curves), as well as the ideal even SC state (dashed curves), for different system parameters. c and d show the corresponding squeezing parameter $r^{\prime}$ of the target squeezed even SC state for which the fidelities in a and b are optimized. The squeezing parameters of the initial SV states are r=0.7r=-0.7 and r=2r=-2 for a and b, respectively. The other parameters are the same as those used in Fig. 2.
@@
Figure 7a shows the Wigner function of $|\Psi\rangle$ for r=0.7r=-0.7, while Fig. 7b displays the Wigner function of the squeezed even SC state with $r^{\prime}=-0.14$ and $\alpha=2$. For these parameters, the fidelity between the two states is approximately $F_{2}\approx 0.97$. This further confirms the high similarity of our generated output signal state $|\Psi\rangle$ to the $|\Psi_{sscs}\rangle$ with large amplitude $\alpha$. Fig. 7 also illustrates the effects of our protocol on the initially prepared Gaussian SV state, which transforms it into a nG state.
Figure 7: Wigner functions of the generated state $|\Psi\rangle$. a Wigner function of the generated state $|\Psi\rangle$ with r=0.7r=-0.7. b Wigner function of the ideal squeezed even SC state with $r^{\prime}=-0.14$ and $\alpha=2$. Other parameters are the same as those used in Fig. 2. For these parameters, the fidelity calculated using the Wigner functions of the generated and ideal squeezed even SC states is $F_{2}\approx 0.97$.
As mentioned previously, the generation of SC states with high fidelity ($F\geq 0.99$) while maintaining large coherent amplitudes ($\alpha\geq 2$) is a crucial task. Although our scheme does not directly generate large-amplitude SC states, the output signal state achieves exceptionally high fidelity with a squeezed even SC state possessing a large amplitude $\alpha$. Therefore, our approach could be applied to problems where large-amplitude SC states are required, offering potentially better performance.
Enlargement of Schrödinger Cat States
@@
where $\kappa^{\prime}=g\langle n_{a}\rangle_{w}$ and $\mathcal{N}_{\pm}^{\prime}$ is the normalization constant, given by
@@
The success probability for this process is given by $\epsilon^{2}|\mathcal{N_{\pm}^{\prime}}|^{-2}$. For the parameters considered here ($\epsilon\thicksim 10^{-3}$, $g=0.01$), this results in probabilities in the range of $10^{-6}$ to $10^{-3}$, with higher fidelities typically corresponding to the more favorable end of this range.
@@
Figure 8: Optimal fidelities and enlargement coefficients. a and c show the optimal fidelities between the output signal states $|\Psi_{\pm}\rangle$ and the ideal SC states $|SCS_{\pm}(\alpha^{\prime})\rangle$, plotted as a function of the coherent state amplitude $\alpha$. b and d show the corresponding enlargement coefficients $\eta_{-}=\alpha^{\prime}/\alpha$ and $\eta_{+}=\alpha^{\prime}/\alpha$, for the odd and even SC states, respectively, for which the fidelities in a and c are optimized. The other parameters are the same as those used in Fig. 2.
In Fig. 8, we show the optimal fidelity functions $F_{+}$ and $F_{-}$ between our generated output states $|\Psi_{+}\rangle$ and $|\Psi_{-}\rangle$ and the ideal even and odd SC states, respectively, as functions of coherent state amplitude $\alpha$. The corresponding optimized amplitudes $\alpha^{\prime}$ of the ideal SC states are used for the maximal overlap. As shown in Fig. 8a, the generated state $|\Psi_{-}\rangle$ provides an excellent approximation to the ideal odd SC state, with fidelity $F_{-}\geq 0.965$ across a wide range of amplitudes, $\alpha$. Notably, in the regions $0<\alpha<0.8$ and $\alpha\geq 2.45$, the fidelity exceeds 0.99. For the even SC state input case [Fig. 8c], the generated state $|\Psi_{+}\rangle$ initially shows noticeable deviation from the ideal even SC state in the small-amplitude regime ($0<\alpha<0.29$). However, as $\alpha$ increases beyond 0.29, the fidelity $F_{+}$ improves significantly, surpassing $0.9$ for all $\alpha>1.24$ regions. Thus, for $\alpha>1.24$, $|\Psi_{+}\rangle$ also can be considered an excellent approximation to the ideal even SC state. Moreover, in high-fidelity regions, we observed that the optimized coherent state amplitudes $\alpha^{\prime}$ of the ideal SC states are consistently larger than the corresponding values of $\alpha$ for our generated states $|\Psi_{\pm}\rangle$ [see Figs. 8b and 8d]. This indicates that our protocol effectively enlarges the SC states. This enlargement effect can be further confirmed by comparing the Wigner functions of the input ideal SC states and the generated output states $|\Psi_{\pm}\rangle$.
@@
[…]